OpenStudy (anonymous):

Help! Determine the x-intercept of the function f(x) = 2^x - 4

OpenStudy (anonymous):

\[2^{x-4}\] or \[2^x-4\]?

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

solve \[2^x-4=0\] in your head pretty much

OpenStudy (anonymous):

clear or no?

OpenStudy (isaidavila):

2^x-4=0 2^x=4 2^x=2^2 x=2

OpenStudy (anonymous):

yes thank you!